OpenStudy (anonymous):

Simplify b^3·b^6. Answer (A) b^2 (b) B^3 (c) B^9 (d)B^18

OpenStudy (anonymous):

b^3.b^6=b^(6+3) b^9 = powers of same number adds when we multiply both. . .

OpenStudy (anonymous):

B: B^9 here's the law of exponent for your reference in the future

OpenStudy (anonymous):

\[\large \color{green}{b^a \cdot b^b = b^{a+b}}\]

OpenStudy (anonymous):

http://www.purplemath.com/modules/exponent.htm

OpenStudy (anonymous):

@huneya they will add up not multiply..